OpenStudy (derpisthewordorisit):

Find the solution set for the given inequality if the replacement set is {1, 2, 3, 4, 5, 6}. 3x – 5 ≥ 7 A. {1, 2, 3, 4} B. {1, 2, 3} C. {4, 5, 6} D. {5, 6} @forgiveness

OpenStudy (anonymous):

Substitute all the numbers in the answers

OpenStudy (derpisthewordorisit):

Okay, this may take a while, I'll check them all.

OpenStudy (anonymous):

When you substitute 1, you end up with: 3 x 1 - 5 3 - 5 -2

OpenStudy (anonymous):

So 1 doesn't work

OpenStudy (derpisthewordorisit):

So only C and D

OpenStudy (anonymous):

Yes, but C gives more data.

OpenStudy (anonymous):

C would be the better answer. Good Job!

OpenStudy (derpisthewordorisit):

Okay! Just 2 more and we're all done.

OpenStudy (anonymous):

\[3x-5\ge7\] \[3x \ge7+5\] \[x \ge \frac{ 12 }{ 3 }\] \[x \ge 4\] solution is ?